VEGAS Pulsar Project Meeting: 2015 July 13th 2:00 - 3:00pm ET

Room / Connection Details

  • GB-137 / Soc 280
  • 192.33.117.12##7144
  • 434-817-6443

Housekeeping Details

  • Main wiki page
  • gbsapp Mailing List
  • Location and regular date and time of meeting - agreed weekly on Mondays; GB-137 / Soc-280; 2-3pm ET
  • ETK codes:
    • Electronics: 432510.GB3015
    • Software: 893252.GB3015
    • Science: 432540.GB3015

Resource Allocations

GUPITA, ARINDAM 1.00
BRANDT, JOE 0.1
FORD, JOHN 0.05
MCCULLOUGH, RANDY 0.05
RAY, JASON 0.55
BLOSS, MARTY 0.1
CREAGER, RAY 0.65
RICHMOND DECKER, JUSTIN 0.2
LYNCH, RYAN 0.25
DEMOREST, PAUL 0.1
PRESTAGE, RICHARD 0.05

Present

  • Arindam, Richard, Ray, Marty, Ryan, Randy, Paul, Scott

Agenda

General Updates

  • BOF Files
    • Incoherent mode builds continuing as planned
    • c1500x0512 mode built without timing errors, but had to manually optimize. Still working on other coherent modes.
  • Manager/software
    • Ray compiling a net task for new HPC pgragram
    • Ryan seeing errors with bof file tests on south and srbs-hpc1 * After some seemingly random number of scans the dealer exits due to a timeout in waiting for NETSTAT to switch to receiving * Appears with new and old bofs * On at least one occasion the shared memory crashed but this doesn't seem to happen every time * Ray and Ryan still investigating
  • GBT Tests
    • Used C-band observing only the noise diode
    • GBT configured for spectral line scans through Astrid, then VEGAS re-configured for pulsar modes using grail client
    • Config tool threw some errors due to the new VEGAS roach device names, but Justin and Ray worked around this
    • Shared memory crashed a few times will running status with the manager; recommended to not continually run status for now
    • Were able to record psrfits data from multiple banks but we did not see the expected bandpass, and in some cases got no signal
      • The IF manager was reporting some unexpected values for bandwidth, sky frequency, and IF frequency, but a manual inspection of each component in cleo seemed to show the correct set-up
      • Ryan was able to successfully configure beef and take a cal scan
      • Need to double check that IF path is being configured properly; will try to take spectral line scans
    • Still planning on observing a pulsar by end of next week, if possible.
  • Wideband AP
    • Circuit design plotted
    • Randy hopes to have things ready to hand off to an engineer by end of next week

Date / Time of Next Meeting

  • July 20

-- RyanLynch - 2015-07-15
Topic revision: r1 - 2015-07-16, RyanLynch
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback